OpenStudy (anonymous):

The sum of three consecutive numbers is 84. What is the largest of these numbers?= 26?

jimthompson5910 (jim_thompson5910):

If the largest is 26, then the 3 numbers are 24,25,26 add them: 24+25+26 = 75 which is not 84 like we want

jimthompson5910 (jim_thompson5910):

x = first number x+1 = second number x+2 = third number add them up and you force that sum to be 84 x + (x+1) + (x+2) = 84 solve for x and then compute x+2 to get your answer

OpenStudy (anonymous):

26,27,28,29 are the choices

OpenStudy (anonymous):

3x=84 x=28

OpenStudy (anonymous):

so i would be 26...28...30 smallest would be 26?

jimthompson5910 (jim_thompson5910):

you should have 3x+3 = 84

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

yes, x = 27 making x+2 = 29

OpenStudy (anonymous):

largest is 29

OpenStudy (anonymous):

29-2=27?

jimthompson5910 (jim_thompson5910):

yeah so the answer is 29

jimthompson5910 (jim_thompson5910):

27+28+29 = 84

OpenStudy (anonymous):

thanks you much